OpenStudy (anonymous):

Simplify each radical expression. Leave in radical form. Show your work. sqrt 45n^5

OpenStudy (cggurumanjunath):



OpenStudy (amorfide):

\[\sqrt{45n^{5}}\] to simplify a root, you want to find factors of 45, which one of them must be a square number you want the largest square number you can get that goes into it so you know that the largest square number that goes into 45 is 9 9 multiplied by 5 goes into 45 therefore you re write 45 as \[\sqrt{9 \times 5n^{5}}\] \[\sqrt{a \times b}=\sqrt{a}\sqrt{b}\] to get \[\sqrt{9}\sqrt{5n^{5}}\]

OpenStudy (cggurumanjunath):